# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v2

@@ Abstract
This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions. First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs. Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results. Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive. Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
@@
Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks. The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process. This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance).
Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020). Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al. (2022b), Section 3.1 in Ma et al. (2022), Section 6 in Zhang et al. (2022) to state a few).
@@
  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs. Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs.
  • Solution to the LASSO problem: We show that NGDMs do not produce sparse solutions for the LASSO problem even for the simplest case of the $L_{1}$ penalized linear model. Furthermore, we show that the NGDM solution to the LASSO problem can produce a counterintutive result – increasing the LASSO penalty can result in a solution with a larger $L_{1}$ norm.
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to Ma et al. (2022)’s claim that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
@@ 1 Preliminaries
In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics. Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis. Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014). This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions. The present section addresses background and terminology that is pertinent to all the subsequent sections.
@@
In this paper, we show the problems with this assumption using a variety of convex loss functions. The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex. The loss function for this problem is given by
@@
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, and $\lambda_{2}\geq 0$ is the ridge penalty. Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations is $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
@@
Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of Boyd et al. (2003) and Tibshirani (2015). The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions. We reference the relevant portions from the lecture notes as they arise.
@@
Our formulations in equations (1) and (2) employ a constant learning rate. This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details). In the GD literature, the training regimen described in equation (1) is called a constant step-size regime. In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing. Although constant step-size GDs approach the stationary point, they rarely converge to it except in special cases.
To ensure provable convergence to the stationary point, theoretical studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
@@
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit.
In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training. The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime. In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution. Essentially, the reducing step-size regime involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. In the reducing step size regime, the theoretical guarantees described in the diminishing step-size regime may not be realized. Nevertheless, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run. Accordingly, in this paper, we exclusively study the constant step-size problem, assuming that the weights and learning rates have been appropriately chosen.
@@ 2 NDGM and convergence analysis
Recent theoretical work by Davis et al. (2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. Unfortunately, these regularity conditions are rarely satisfied in practice. To illustrate this point, consider the Assumption A.3 in Davis et al. (2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge. As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs. However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s). With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
@@
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$. Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$. Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$. As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction. The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of Boyd et al. (2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
@@
Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training. Majority of the analyses in the literature (Défossez et al. (2020); Reddi et al. (2019); You et al. (2019); Zhang et al. (2022) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
@@
for all $x$, $y$ in the domain of $f$. However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain. This leads us to a pertinent inquiry: Can we salvage this assumption by confining its validity to a local context, possibly within a specific interval centered around the local optimal solution?
@@
would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) likely have little bearing on the dynamics of convergence in modern non-smooth neural networks. More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
@@
To validate our intuition in a more general context, we conduct experiments on CNNs of various depths (see the appendix for architecture details). For each network, we carry out two identical simulations, with the only difference being the activation function used. In the first simulation, we employ the non-smooth ReLU activation function, and in the second, we use the similarly shaped, but smooth and asymptotally equivalent, GELU activation function (Hendrycks & Gimpel, 2016).
As seen in Figure 2, the difference in convergence speeds between smooth and non-smooth networks widens as the depth of the neural network. This observation aligns with our expectations, given that increased depth introduces greater non-smoothness into the system. Hendrycks & Gimpel (2016) substantiate a similar trend of accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19 respectively. As the network depth increases, the non-smoothness in the system also increases. Consequently, the difference in convergence speeds between smooth and non-smooth networks widens with the increasing non-smoothness in the system.
Do results derived in the L-Smooth setting describe the behavior of all differentiable neural networks?:
In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks. We point out that $L-smoothness$ is one of the strongest assumptions which can be applied to a function after convexity. This assumption guarantees the existence of the first and second derivatives, and the boundedness of the second derivative—a collection of criteria that are unmet by the majority of the loss functions encountered in the neural network literature.
@@
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) which employs the differentiable GELU activation function (Hendrycks & Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.

### v3

@@ Abstract
This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions. First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs. Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results. Additionally, we dispel the common belief that optimization algorithms like Adam and RMSProp perform similarly in non-differentiable contexts. Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive. Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
@@
Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks. The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process. This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance). To emphasize this point, we note that three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not mention non-differentiability, even though the ReLU, arguably the most common activation function in neural networks, is non-differentiable.
Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes ((Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019) to state a few), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020). Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in (Ahn et al., 2022b), Section 3.1 in (Ma et al., 2022), Section 6 in (Zhang et al., 2022) to state a few).
@@
  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs. Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs. Furthermore, we dispel the widespread belief that different NDGM variants like RMSProp and Adam largely differ only in their rates of convergence, but converge to roughly the same eventual point in the limit.
  • Solution to the LASSO problem: We show that NGDMs, including variants like RMSProp and NDGM with momentum, do not produce sparse solutions for the LASSO problem, even for the simplest case of the $L_{1}$ penalized linear model, and exhibit different behaviors, contrary to prevalent wisdom. Moreover, NGDMs can yield counterintuitive results, such as producing a solution with a larger $L_{1}$ norm with a larger LASSO penalty.
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to the claim made in (Ma et al., 2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
@@ 1 Preliminaries
In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics. Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis. Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014). This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions. Our experiments have relatively low computational costs, with the most expensive run completing within 2 hours on a single GPU. The present section addresses background and terminology that is pertinent to all the subsequent sections.
@@
In this paper, we show the problems with this assumption using a variety of convex loss functions. The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which (Kumar, 2023) has shown to be convex with with the all-zeros vector as its unique global minimizer. The loss function for this problem is given by
@@
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$. Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
@@
Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of (Boyd et al., 2003) and (Tibshirani, 2015). The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions. We reference the relevant portions from the lecture notes as they arise.
@@
Our formulations in equations (1) and (2) employ a constant learning rate. This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details). In the GD literature, the training regimen described in equation (1) is called a constant step-size regime. In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing. Although constant step-size GDs approach the stationary point, they rarely converge to it unless the loss function possesses desirable properties like $L-smoothness$ (Proposition 1.2.3 in (Bertsekas, 1997)). To ensure provable convergence to the stationary point in more general settings, studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
@@
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit. In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training. The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime.
In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution. In the reducing step-size regime, training involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. While the theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run. Accordingly, we exclusively study the constant step-size problem, assuming appropriate weights and learning rates have been chosen.
@@ 2 NDGM and convergence analysis
Recent theoretical work by (Davis et al., 2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. Unfortunately, the conditions of the diminishing step size regime are rarely satisfied in practice. To illustrate this point, consider the Assumption A.3 in (Davis et al., 2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge. As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs. However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s). With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
@@
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$. Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$. Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$. As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction. The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of (Boyd et al., 2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
@@
Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training. Majority of the analyses in the literature ((Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
@@
for all $x$, $y$ in the domain of $f$. However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain. This raises the question if the assumption can be salvaged by confining its validity to a local context around the local optimal solution.
@@
would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of (Tibshirani, 2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) have little bearing on the dynamics of convergence in modern non-smooth neural networks. More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
@@
We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A for architecture details). In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks & Gimpel, 2016). The first set of experiments excludes batch normalization, while the second includes it.
Figure 2 shows that without batch normalization, the gap in convergence speeds between smooth and non-smooth networks widens as network depth increases, aligning with our expectations. This trend is further substantiated by (Hendrycks & Gimpel, 2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper). However, with batch normalization, the difference between ReLU and GELU training curves becomes indistinguishable, emphasizing its importance in the training process. We conclude that the choice of activation function is the primary factor contributing to the difference in convergence speeds between ReLU and GELU curves with increasing depth, as no other principled reason appears evident.
Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19, respectively. The top row shows validation losses without batch normalization, while the bottom row includes batch normalization.
L-smoothness and the geometry of ReLU networks:
As stated in (Bottou, 2010), the assumption of $L-smoothness$ ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector. However, this assumption which does not hold for ReLU networks. To illustrate this, consider the 1-D case with $q(x)=\text{max}(0,x)$. In this case, $\lim_{h\rightarrow 0^{+}}(1/2h)(\nabla q(h)-\nabla q(-h))=\infty$, indicating that the gradient changes arbitrarily quickly at the origin. This observation necessitates the use of non-smooth analysis for studying the loss functions of ReLU networks, as detailed in (Laurent & Brecht, 2018). The key intuition about non-differentiable functions is that they have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks. As a result, gradient descents which rely on smoothness assumptions cannot access these points of interest. This realization serves as one of the primary motivations for the development and application of second-order cone programs for convex non-smooth problems (Boyd et al., 2004)  [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh. The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme throughout the remaining sections of this paper. In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks. $L-smoothness$ is one of the strongest assumptions after convexity, guaranteeing the existence and boundedness of the first and second derivatives—criteria unmet by most loss functions in neural network literature.
Limitations of our work:
A common limitation in the current and subsequent sections is that we do not analyze the dynamics of the neural network training process as it is often conducted in practice. This is necessitated by the fact that neural networks often involve non-differentiable layers like dropout and max pooling, which are beyond the scope of our study. We do not reiterate this limitation in every subsequent section.
@@
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks & Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.